# Enhancing Conversational Agents from Open-Source Large Language Models with Illocutionary Force and Document-Based Knowledge Retrieval

### v1

@@ 0 $\Rightarrow$ indicates a poor performance
Without Illocutionary Force With Illocutionary Force
@@
    Response Time: It is observed that the response time of the chat agent was faster in all the models for the experiment which included the illocutionary force except for the llama2:13b model in which it was faster in the experiment without the illocutionary force included in the user query.
@@
    BERT-Precision: It is observed that the semantic similarities based on the BERT-embeddings were better in Llama3-chatqa-latest and Llama3-latest models in the experiment without including the illocutionary force, and was better in the rest of the other models in the experiment which included the illocutionary force.
@@
    BERT-Recall: It is observed that the breadth and coverage similarities based on the BERT-embeddings of the chat agent response and the custom document were better in Llama3-chatqa-latest and Llama3-latest models in the experiment without including the illocutionary force, and was better in the rest of the other models in the experiment which included the illocutionary force.
@@
    BERT-F1: It is observed that a balanced trade-off between the precision and recall which gives a better semantic balance based on the BERT-embeddings of the chat agent response and the custom document were better in Llama3-chatqa-latest and Llama3-latest models in the experiment without including the illocutionary force, and was better in the rest of the other models in the experiment which included the illocutionary force.
@@
    QA-Ref: Comparing the chat agent response typically with answers in the knowledge document, it is observed that all the models in both the experiment without and with the illocutionary force included had the same performance indicating a balanced metric.
@@
    QA-Cand: When we observed the flexibility of the response with respect to the custom document provided, it was seen that llama2:13b, Tinyllama:latest, Llama3-chatqa-latest performed better in the experiment with the illocutionary force included, performed same with llama3-latest, and performed better in mistral:latest for the experiment without the illocutionary force.
@@
    ROUGE-1: It is observed that elements of unigram keywords in the chat agent response found in the custom document were better in llama2:13b and Llama3-chatqa-lates models in the experiment without including the illocutionary force, and were better in the rest of the other models in the experiment which included the illocutionary force.
@@
    ROUGE-2: It is observed that elements of Bigram indicating coherence in the chat agent response based on the custom document were better in llama2:13b and Llama3-chatqa-lates models in the experiment without including the illocutionary force, and were better in the rest of the other models in the experiment which included the illocutionary force.
@@
    ROUGE-L: It is observed that the longest common subsequence between the chat agent response and the custom document was better in llama2:13b and Llama3-chatqa-latest models in the experiment without including the illocutionary force, and was better in the rest of the other models in the experiment which included the illocutionary force.
@@
    METEOR: It is observed that a better linguistically diverse yet semantically response from the chat agent with respect to the custom document, was better in llama2:13b and Llama3-chatqa-latest models for the experiment conducted without including the illocutionary force, and better in the other models for the experiment conducted with the inclusion of the illocutionary force with the user query.
@@
    Perplexity: The ability to predict the following words in a sequence scenario was better in all the models for the experiment conducted without including the illocutionary force with a lower perplexity value as compared to same models for the experiment conducted with the inclusion of the illocutionary force with the user query.
@@
    Avg Total: Taking a record of the performance of the models across all metrics and excluding both the response time and the results of the similar values, it is observed that: the llama2:13b model performs better with an average score of 5 in the experiment without the illocutionary force against same model for the experiment with the illocutionary force which had a score of 4; the Tinyllama:latest model performs better with an average score of 8 in the experiment with the illocutionary force against same model for the experiment without the illocutionary force which had a score of 1; the Llama3-chatqa-latest model performs better with an average score of 8 in the experiment without the illocutionary force against same model for the experiment with the illocutionary force which had a score of 1; the Llama3-latest model had similar score of 4 in the experiment with the illocutionary force and for the experiment without the illocutionary force which also had a score of 4; the Mistral:latest model performs better with an average score of 7 in the experiment with the illocutionary force against same model for the experiment without the illocutionary force which had a score of 2.
@@
Without Illocutionary Force With Illocutionary Force
@@
    Response Time: It is observed that the response time of the chat agent was faster in Llama3-latest and Mistral:latest for the experiment which included the illocutionary force and was faster for the other models in the experiment without the illocutionary force included in the user query.
@@
    BERT-Precision: It is observed that the semantic similarities based on the BERT-embeddings were better in llama2:13b and Tinyllama:latest models in the experiment without including the illocutionary force, and were better in the rest of the other models in the experiment which included the illocutionary force.
@@
    BERT-Recall: It is observed that the breadth and coverage similarities based on the BERT-embeddings of the chat agent response and the custom document were better in llama2:13b and Tinyllama:latest models in the experiment without including the illocutionary force, and was better in the rest of the other models in the experiment which included the illocutionary force.
@@
    BERT-F1: It is observed that a balanced trade-off between the precision and recall which gives a better semantic balance based on the BERT-embeddings of the chat agent response and the custom document were better in llama2:13b and Tinyllama:latest models in the experiment without including the illocutionary force, and was better in the rest of the other models in the experiment which included the illocutionary force.
@@
    QA-Ref: Comparing the chat agent response typically with answers in the knowledge document, it is observed that Tinyllama:latest, Llama3-latest, and Mistral:latest models in both the experiment without and with the illocutionary force included had the same performance indicating a balanced metric, but was better llama2:13b, and Llama3-chatqa-latest in the experiment with the inclusion of the illocutionary force as compared to the experiment without the illocutionary force.
@@
    QA-Cand: When we observed the flexibility of the response with respect to the custom document provided, it was seen that Llama3-chatqa-latest, and Mistral:latest performed better in the experiment with the illocutionary force included, and performed better in other models for the experiment without the illocutionary force.
@@
    ROUGE-1: It is observed that elements of unigram keywords in the chat agent response found in the custom document were better in llama2:13b, Tinyllama:latest, and Llama3-chatqa-latest models in the experiment without including the illocutionary force, and were better in the rest of the other models in the experiment which included the illocutionary force.
@@
    ROUGE-2: It is observed that elements of bigram indicating coherence in the chat agent response based on the custom document were better in all the models in the experiment without including the illocutionary force, and was same in only the Mistral:latest model in the experiment which included the illocutionary force.
@@
    ROUGE-L: It is observed that the longest common subsequence between the chat agent response and on the custom document was better in llama2:13b, Tinyllama:latest, and Llama3-chatqa-latest models in the experiment without including the illocutionary force, and were better in the rest of the other models in the experiment which included the illocutionary force.
@@
    METEOR: It is observed the a better linguistically diverse but yet sementically response from the chat agent with respect to the custom document, was better in llama2:13b, Tinyllama:latest, and Llama3-chatqa-latest models for the experiment conducted without including the illocutionary force, and better in the other models for the experiment conducted with the inclusion of the illocutionary force with the user query.
@@
    Perplexity: The ability to predict the following words in a sequence scenario was better in llama2:13b, Llama3-chatqa-latest, and Mistral:latest models for the experiment conducted without including the illocutionary force with a lower perplexity value and better in other models for the experiment conducted with the inclusion of the illocutionary force with the user query.
@@
    Avg Total: Taking a record of the performance of the models across all metrics and excluding both the response time and the results of the similar values, it is observed that: the llama2:13b model performs better with an average score of 9 in the experiment without the illocutionary force against same model for the experiment with the illocutionary force which had a score of 1; the Tinyllama:latest model performs better with an average score of 8 in the experiment without the illocutionary force against same model for the experiment with the illocutionary force which had a score of 1; the Llama3-chatqa-latest model performs same with an average score of 5 in the experiment without the illocutionary force against same model for the experiment with the illocutionary force which had a similar score of 5; the Llama3-latest model performed better with a score of 7 in the experiment with the illocutionary force and for the experiment without the illocutionary force with a score of 2; the Mistral:latest model performs better with an average score of 7 in the experiment with the illocutionary force against same model for the experiment without the illocutionary force which had a score of 1.
Analysis of the two results based on first bot responses in Document 004
Without Illocutionary Force With Illocutionary Force
@@
    Response Time: It is observed that the response time for Llama2:13b, Tinyllama:latest Llama3-latest, and Mistral:latest is faster in the experiment conducted without the illocutionary force as compared with the same models in the experiment with the illocutionary force.
@@
    BERT-Precision: It is observed that the precision metric based on BERT-embeddings showed a better score for Tinyllama:latest and Mistral:latest in the experiment conducted without including the illocutionary force and a better score was observed in the Llama3-latest, and Llama2:13b models for the experiment that included the illocutionary force in the user query.
@@
    BERT-Recall: It is observed that the recall metric based on BERT-embeddings is similar to the precision results with better scores noted for Tinyllama:latest and Mistral:latest in the experiment without including the illocutionary force, and better scores for Llama3-latest and Llama2:13b in the experiment with the illocutionary force included with the user query.
@@
    BERT-F1: It is observed that the balanced precision and recall showed similar results to BERT-Recall and Precision, indicating a better score for Tinyllama:latest and Mistral:latest in the experiment without illocutionary force, and better scores for Llama3-latest and Llama2:13b with the illocutionary force included with the user query.
@@
    QA-Cand: When we observed the flexibility of the response with respect to the custom document provided, it was seen that Llama3-latest and Mistral:latest showed better results in the experiment without the illocutionary force, and better results for Llama2:13b and Tinyllama:latest in the experiment when the illocutionary force was included with the user query.
@@
    ROUGE-1: It is observed that elements of unigram keywords in the chat agent response found in the custom document were better for Mistral:latest in the experiment without the illocutionary force, while Llama2:13b and Tinyllama:latest models showed better scores in the experiment with the illocutionary force included, and a similar score for llama3-latest in both experiments.
@@
    ROUGE-2: It is observed that bigram coherence was better for Llama3-latest and Mistral:latest in the experiment without the illocutionary force, but was observed as similar for Llama2:13b across both experiments and better in Tinyllama:latest with the inclusion of the illocutionary force.
@@
    ROUGE-L: It is observed that the longest common subsequence metric was better for Llama2:13b, Llama3-latest, and Mistral:latest in the experiment conducted without the illocutionary force. However, Tinyllama:latest was the only model to achieve a better result in the experiment with the illocutionary force included in the user query.
@@
    METEOR: It is observed that linguistic diversity and semantic accuracy in responses showed a better score for Llama2:13b, Tinyllama:latest, and Mistral:latest in the experiment with the illocutionary force, but was observed to be similar for Llama3-latest in both experiments.
@@
    Perplexity: It is observed that the predictive ability of the models indicated better values for Tinyllama:latest, Llama3-latest and Mistral:latest in the experiment without the illocutionary force, while the same metric achieved a better score in Llama2:13b for the experiment with the illocutionary force.
@@
    Avg Total: Taking a record of the performance of the models across all metrics, it was observed that the Mistral:latest model performed best with a total score of 8 in the experiment without the illocutionary force and a much lower score of 1 in the experiment with the illocutionary force. Similarly, Tinyllama:latest obtained a score of 4 in the experiment without the illocutionary force but increased to 5 when the illocutionary force was included. Llama2:13b had a significant difference in performance, with a score of 1 in the experiment without illocutionary force but improved to 7 when it was included in the user query. Finally, Llama3-latest showed moderate performance, scoring 4 without the illocutionary force and 3 when it was included in the user query.
Analysis of the two results based on second bot responses in Document 004
Without Illocutionary Force With Illocutionary Force
@@
    Response Time: It is observed that the response time for Llama2:13b, Tinyllama:latest Llama3-latest, and Mistral:latest is faster in the experiment conducted without the illocutionary force as compared with the same models in the experiment with the illocutionary force.
@@
    BERT-Precision: It is observed that the precision metric based on BERT-embeddings showed a better score for llama2:13b, Tinyllama:latest and Llama3-latest in the experiment conducted without including the illocutionary force and a better score was observed in Mistral:latest model for the experiment that included the illocutionary force in the user query.
@@
    BERT-Recall: It is observed that the recall metric based on BERT-embeddings is similar to the precision results with better scores for llama2:13b, Tinyllama:latest and Llama3-latest in the experiment conducted without including the illocutionary force and a better score was observed in Mistral:latest model for the experiment that included the illocutionary force in the user query.
@@
    BERT-F1: It is observed that the balanced precision and recall showed similar results to BERT-Recall and Precision, indicating better scores for llama2:13b, Tinyllama:latest and Llama3-latest in the experiment conducted without including the illocutionary force and a better score was observed in Mistral:latest model for the experiment that included the illocutionary force in the user query.
@@
    QA-Cand: When we observed the flexibility of the response with respect to the custom document provided, it was seen that Llama2:13b, Llama3-latest and Mistral:latest showed better results in the experiment without the illocutionary force, and the same result for Tinyllama:latest in both experiments.
@@
    ROUGE-1: It is observed that elements of unigram keywords in the chat agent response found in the custom document were better for Tinyllama:latest, and llama3-latest in the experiment without the illocutionary force, while Mistral:latest model showed better scores in the experiment with the illocutionary force included, and a similar score for Llama2:13b in both experiments.
@@
    ROUGE-2: It is observed that bigram coherence was better for Llama2:13b and Llama3-latest in the experiment without the illocutionary force, but was observed as similar for Tinyllama:latest across both experiments and better in Mistral:latest with the inclusion of the illocutionary force.
@@
    ROUGE-L: It is observed that the longest common subsequence metric was better for Llama2:13b, Llama3-latest, and Tinyllama:lates in the experiment conducted without the illocutionary force. However, Mistral:latest was the only model to achieve a better result in the experiment with the illocutionary force included in the user query.
@@
    METEOR: It is observed that linguistic diversity and semantic accuracy in responses showed a better score for Llama2:13b, and Llama3-latest in the experiment with the illocutionary force, but was observed to be better for Tinyllama:latest, and Mistral:latest in the experiment with the illocutionary force included in the user query.
@@
    Perplexity: It is observed that the predictive ability of the models indicated better values for Llama2:13b, Llama3-latest and Mistral:latest in the experiment without the illocutionary force, while the same metric achieved a better score in Tinyllama:latest for the experiment with the illocutionary force.
@@
    Avg Total: Taking a record of the performance of the models across all metrics, it was observed that the Llama3-latest model performed best with a total score of 10 in the experiment without the illocutionary force and a much lower score of 0 in the experiment with the illocutionary force. Similarly, llama2:13b obtained a score of 8 in the experiment without the illocutionary force but dropped to 0 when the illocutionary force was included. Tinyllama:latest showed a difference in performance, with a score of 5 in the experiment without illocutionary force but dropped to 3 when it was included in the user query. Finally, Mistral:latest showed a performance of 2 without the illocutionary force and was improved to 7 when it was included in the user query.

### v2

@@ 0 $\Rightarrow$ indicates a poor performance
Without speech acts With speech acts
@@
    Response Time: It is observed that the response time of the chat agent was faster in all the models for the experiment which included the speech acts except for the llama2:13b model in which it was faster in the experiment without the speech acts included in the user query.
@@
    BERT-Precision: It is observed that the semantic similarities based on the BERT-embeddings were better in Llama3-chatqa-latest and Llama3-latest models in the experiment without including the speech acts, and was better in the rest of the other models in the experiment which included the speech acts.
@@
    BERT-Recall: It is observed that the breadth and coverage similarities based on the BERT-embeddings of the chat agent response and the custom document were better in Llama3-chatqa-latest and Llama3-latest models in the experiment without including the speech acts, and was better in the rest of the other models in the experiment which included the speech acts.
@@
    BERT-F1: It is observed that a balanced trade-off between the precision and recall which gives a better semantic balance based on the BERT-embeddings of the chat agent response and the custom document were better in Llama3-chatqa-latest and Llama3-latest models in the experiment without including the speech acts, and was better in the rest of the other models in the experiment which included the speech acts.
@@
    QA-Ref: Comparing the chat agent response typically with answers in the knowledge document, it is observed that all the models in both the experiment without and with the speech acts included had the same performance indicating a balanced metric.
@@
    QA-Cand: When we observed the flexibility of the response with respect to the custom document provided, it was seen that llama2:13b, Tinyllama:latest, Llama3-chatqa-latest performed better in the experiment with the speech acts included, performed same with llama3-latest, and performed better in mistral:latest for the experiment without the speech acts.
@@
    ROUGE-1: It is observed that elements of unigram keywords in the chat agent response found in the custom document were better in llama2:13b and Llama3-chatqa-lates models in the experiment without including the speech acts, and were better in the rest of the other models in the experiment which included the speech acts.
@@
    ROUGE-2: It is observed that elements of Bigram indicating coherence in the chat agent response based on the custom document were better in llama2:13b and Llama3-chatqa-lates models in the experiment without including the speech acts, and were better in the rest of the other models in the experiment which included the speech acts.
@@
    ROUGE-L: It is observed that the longest common subsequence between the chat agent response and the custom document was better in llama2:13b and Llama3-chatqa-latest models in the experiment without including the speech acts, and was better in the rest of the other models in the experiment which included the speech acts.
@@
    METEOR: It is observed that a better linguistically diverse yet semantically response from the chat agent with respect to the custom document, was better in llama2:13b and Llama3-chatqa-latest models for the experiment conducted without including the speech acts, and better in the other models for the experiment conducted with the inclusion of the speech acts with the user query.
@@
    Perplexity: The ability to predict the following words in a sequence scenario was better in all the models for the experiment conducted without including the speech acts with a lower perplexity value as compared to same models for the experiment conducted with the inclusion of the speech acts with the user query.
@@
    Avg Total: Taking a record of the performance of the models across all metrics and excluding both the response time and the results of the similar values, it is observed that: the llama2:13b model performs better with an average score of 5 in the experiment without the speech acts against same model for the experiment with the speech acts which had a score of 4; the Tinyllama:latest model performs better with an average score of 8 in the experiment with the speech acts against same model for the experiment without the speech acts which had a score of 1; the Llama3-chatqa-latest model performs better with an average score of 8 in the experiment without the speech acts against same model for the experiment with the speech acts which had a score of 1; the Llama3-latest model had similar score of 4 in the experiment with the speech acts and for the experiment without the speech acts which also had a score of 4; the Mistral:latest model performs better with an average score of 7 in the experiment with the speech acts against same model for the experiment without the speech acts which had a score of 2.
@@
Without speech acts With speech acts
@@
    Response Time: It is observed that the response time of the chat agent was faster in Llama3-latest and Mistral:latest for the experiment which included the speech acts and was faster for the other models in the experiment without the speech acts included in the user query.
@@
    BERT-Precision: It is observed that the semantic similarities based on the BERT-embeddings were better in llama2:13b and Tinyllama:latest models in the experiment without including the speech acts, and were better in the rest of the other models in the experiment which included the speech acts.
@@
    BERT-Recall: It is observed that the breadth and coverage similarities based on the BERT-embeddings of the chat agent response and the custom document were better in llama2:13b and Tinyllama:latest models in the experiment without including the speech acts, and was better in the rest of the other models in the experiment which included the speech acts.
@@
    BERT-F1: It is observed that a balanced trade-off between the precision and recall which gives a better semantic balance based on the BERT-embeddings of the chat agent response and the custom document were better in llama2:13b and Tinyllama:latest models in the experiment without including the speech acts, and was better in the rest of the other models in the experiment which included the speech acts.
@@
    QA-Ref: Comparing the chat agent response typically with answers in the knowledge document, it is observed that Tinyllama:latest, Llama3-latest, and Mistral:latest models in both the experiment without and with the speech acts included had the same performance indicating a balanced metric, but was better llama2:13b, and Llama3-chatqa-latest in the experiment with the inclusion of the speech acts as compared to the experiment without the speech acts.
@@
    QA-Cand: When we observed the flexibility of the response with respect to the custom document provided, it was seen that Llama3-chatqa-latest, and Mistral:latest performed better in the experiment with the speech acts included, and performed better in other models for the experiment without the speech acts.
@@
    ROUGE-1: It is observed that elements of unigram keywords in the chat agent response found in the custom document were better in llama2:13b, Tinyllama:latest, and Llama3-chatqa-latest models in the experiment without including the speech acts, and were better in the rest of the other models in the experiment which included the speech acts.
@@
    ROUGE-2: It is observed that elements of bigram indicating coherence in the chat agent response based on the custom document were better in all the models in the experiment without including the speech acts, and was same in only the Mistral:latest model in the experiment which included the speech acts.
@@
    ROUGE-L: It is observed that the longest common subsequence between the chat agent response and on the custom document was better in llama2:13b, Tinyllama:latest, and Llama3-chatqa-latest models in the experiment without including the speech acts, and were better in the rest of the other models in the experiment which included the speech acts.
@@
    METEOR: It is observed the a better linguistically diverse but yet sementically response from the chat agent with respect to the custom document, was better in llama2:13b, Tinyllama:latest, and Llama3-chatqa-latest models for the experiment conducted without including the speech acts, and better in the other models for the experiment conducted with the inclusion of the speech acts with the user query.
@@
    Perplexity: The ability to predict the following words in a sequence scenario was better in llama2:13b, Llama3-chatqa-latest, and Mistral:latest models for the experiment conducted without including the speech acts with a lower perplexity value and better in other models for the experiment conducted with the inclusion of the speech acts with the user query.
@@
    Avg Total: Taking a record of the performance of the models across all metrics and excluding both the response time and the results of the similar values, it is observed that: the llama2:13b model performs better with an average score of 9 in the experiment without the speech acts against same model for the experiment with the speech acts which had a score of 1; the Tinyllama:latest model performs better with an average score of 8 in the experiment without the speech acts against same model for the experiment with the speech acts which had a score of 1; the Llama3-chatqa-latest model performs same with an average score of 5 in the experiment without the speech acts against same model for the experiment with the speech acts which had a similar score of 5; the Llama3-latest model performed better with a score of 7 in the experiment with the speech acts and for the experiment without the speech acts with a score of 2; the Mistral:latest model performs better with an average score of 7 in the experiment with the speech acts against same model for the experiment without the speech acts which had a score of 1.
Analysis of the two results based on first bot responses in Document 002
Without speech acts With speech acts
@@
    Response Time: It is observed that the response time for Llama2:13b, Tinyllama:latest Llama3-latest, and Mistral:latest is faster in the experiment conducted without the speech acts as compared with the same models in the experiment with the speech acts.
@@
    BERT-Precision: It is observed that the precision metric based on BERT-embeddings showed a better score for Tinyllama:latest and Mistral:latest in the experiment conducted without including the speech acts and a better score was observed in the Llama3-latest, and Llama2:13b models for the experiment that included the speech acts in the user query.
@@
    BERT-Recall: It is observed that the recall metric based on BERT-embeddings is similar to the precision results with better scores noted for Tinyllama:latest and Mistral:latest in the experiment without including the speech acts, and better scores for Llama3-latest and Llama2:13b in the experiment with the speech acts included with the user query.
@@
    BERT-F1: It is observed that the balanced precision and recall showed similar results to BERT-Recall and Precision, indicating a better score for Tinyllama:latest and Mistral:latest in the experiment without speech acts, and better scores for Llama3-latest and Llama2:13b with the speech acts included with the user query.
@@
    QA-Cand: When we observed the flexibility of the response with respect to the custom document provided, it was seen that Llama3-latest and Mistral:latest showed better results in the experiment without the speech acts, and better results for Llama2:13b and Tinyllama:latest in the experiment when the speech acts was included with the user query.
@@
    ROUGE-1: It is observed that elements of unigram keywords in the chat agent response found in the custom document were better for Mistral:latest in the experiment without the speech acts, while Llama2:13b and Tinyllama:latest models showed better scores in the experiment with the speech acts included, and a similar score for llama3-latest in both experiments.
@@
    ROUGE-2: It is observed that bigram coherence was better for Llama3-latest and Mistral:latest in the experiment without the speech acts, but was observed as similar for Llama2:13b across both experiments and better in Tinyllama:latest with the inclusion of the speech acts.
@@
    ROUGE-L: It is observed that the longest common subsequence metric was better for Llama2:13b, Llama3-latest, and Mistral:latest in the experiment conducted without the speech acts. However, Tinyllama:latest was the only model to achieve a better result in the experiment with the speech acts included in the user query.
@@
    METEOR: It is observed that linguistic diversity and semantic accuracy in responses showed a better score for Llama2:13b, Tinyllama:latest, and Mistral:latest in the experiment with the speech acts, but was observed to be similar for Llama3-latest in both experiments.
@@
    Perplexity: It is observed that the predictive ability of the models indicated better values for Tinyllama:latest, Llama3-latest and Mistral:latest in the experiment without the speech acts, while the same metric achieved a better score in Llama2:13b for the experiment with the speech acts.
@@
    Avg Total: Taking a record of the performance of the models across all metrics, it was observed that the Mistral:latest model performed best with a total score of 8 in the experiment without the speech acts and a much lower score of 1 in the experiment with the speech acts. Similarly, Tinyllama:latest obtained a score of 4 in the experiment without the speech acts but increased to 5 when the speech acts was included. Llama2:13b had a significant difference in performance, with a score of 1 in the experiment without speech acts but improved to 7 when it was included in the user query. Finally, Llama3-latest showed moderate performance, scoring 4 without the speech acts and 3 when it was included in the user query.
Analysis of the two results based on second bot responses in Document 002
Without speech acts With speech acts
@@
    Response Time: It is observed that the response time for Llama2:13b, Tinyllama:latest Llama3-latest, and Mistral:latest is faster in the experiment conducted without the speech acts as compared with the same models in the experiment with the speech acts.
@@
    BERT-Precision: It is observed that the precision metric based on BERT-embeddings showed a better score for llama2:13b, Tinyllama:latest and Llama3-latest in the experiment conducted without including the speech acts and a better score was observed in Mistral:latest model for the experiment that included the speech acts in the user query.
@@
    BERT-Recall: It is observed that the recall metric based on BERT-embeddings is similar to the precision results with better scores for llama2:13b, Tinyllama:latest and Llama3-latest in the experiment conducted without including the speech acts and a better score was observed in Mistral:latest model for the experiment that included the speech acts in the user query.
@@
    BERT-F1: It is observed that the balanced precision and recall showed similar results to BERT-Recall and Precision, indicating better scores for llama2:13b, Tinyllama:latest and Llama3-latest in the experiment conducted without including the speech acts and a better score was observed in Mistral:latest model for the experiment that included the speech acts in the user query.
@@
    QA-Cand: When we observed the flexibility of the response with respect to the custom document provided, it was seen that Llama2:13b, Llama3-latest and Mistral:latest showed better results in the experiment without the speech acts, and the same result for Tinyllama:latest in both experiments.
@@
    ROUGE-1: It is observed that elements of unigram keywords in the chat agent response found in the custom document were better for Tinyllama:latest, and llama3-latest in the experiment without the speech acts, while Mistral:latest model showed better scores in the experiment with the speech acts included, and a similar score for Llama2:13b in both experiments.
@@
    ROUGE-2: It is observed that bigram coherence was better for Llama2:13b and Llama3-latest in the experiment without the speech acts, but was observed as similar for Tinyllama:latest across both experiments and better in Mistral:latest with the inclusion of the speech acts.
@@
    ROUGE-L: It is observed that the longest common subsequence metric was better for Llama2:13b, Llama3-latest, and Tinyllama:lates in the experiment conducted without the speech acts. However, Mistral:latest was the only model to achieve a better result in the experiment with the speech acts included in the user query.
@@
    METEOR: It is observed that linguistic diversity and semantic accuracy in responses showed a better score for Llama2:13b, and Llama3-latest in the experiment with the speech acts, but was observed to be better for Tinyllama:latest, and Mistral:latest in the experiment with the speech acts included in the user query.
@@
    Perplexity: It is observed that the predictive ability of the models indicated better values for Llama2:13b, Llama3-latest and Mistral:latest in the experiment without the speech acts, while the same metric achieved a better score in Tinyllama:latest for the experiment with the speech acts.
@@
    Avg Total: Taking a record of the performance of the models across all metrics, it was observed that the Llama3-latest model performed best with a total score of 10 in the experiment without the speech acts and a much lower score of 0 in the experiment with the speech acts. Similarly, llama2:13b obtained a score of 8 in the experiment without the speech acts but dropped to 0 when the speech acts was included. Tinyllama:latest showed a difference in performance, with a score of 5 in the experiment without speech acts but dropped to 3 when it was included in the user query. Finally, Mistral:latest showed a performance of 2 without the speech acts and was improved to 7 when it was included in the user query.